Calculate the magnetic field strength needed on a 449-turn square loop 15 cm on a side to create a maximum torque of 322 N⋅m if the loop is carrying 25 A.


answer to 2 decimal places


1
Expert's answer
2021-05-27T09:59:15-0400

Since the magnitude of the torque on a coil of N loops each has the area A, carries a current I and its area vector makes an angle θ with a magnetic field of density B is given by the equation:

"|\vec{τ}|=NIABsinθ"

Therefore, the maximum value of the torque is when θ = 90°, with the magnitude of the torque given by:

"|\vec{τ}_{max}|=BIAN"

From this relation we get that:

"B = \frac{|\vec{τ}_{max}|}{IAN} \\

= \frac{322}{25 \times 0.15^2 \times 449} \\

= 1.27 \;T"